OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1)

OpenStudy (anonymous):

K so we will seperate y in the first equation. 2x+y=3 -2x -2x ---------- y=-2x+3 Now all we do is put this as our y value in our 2nd equation -3x+4(-2x+3)=1 -3x-8x+12=1 [We add like terms] -11x+12=1 [Subtract 12 on both sides] -11x=-11 [Divide both sides by -11] x=1 Now we can put into our 1st equation and solve- 2(1)+y=3 2+y=3 [Subtract both sides by 2] y=1 This means our answer is (1,1) -Hope that helps :)

OpenStudy (anonymous):

It does sir..! Thank you ♥

OpenStudy (anonymous):

No problem :)